# Constrained mean-field control with singular control: Existence, stochastic maximum principle and constrained FBSDE

### v1

@@ 2.2 Standing assumptions
    Let us define
    $F(t,x,\mu,u,y,z):=\nabla_{x}H(t,x,\mu,u,y,z)+\int_{\mathbb{R}^{n}}\partial_{\mu}H(t,x^{\prime},\mu,u,y,z)(x)\mu(\mathrm{d}x^{\prime}),$
    and set
    $\hat{F}(t,x,\mu,y,z):=F(t,x,\mu,\hat{u}(t,x,\mu,y,z),y,z),$
    $\Phi(t,x,\mu,y,z):=(-\hat{F},\hat{b},\hat{\sigma})(t,x,\mu,y,z).$
    The following monotone condition holds that
@@
    where $\beta>\frac{M}{2}$ with $M$ being given in (A3) and the inner product is read as
@@
    $\hat{b}$, $\hat{\sigma}$, $\hat{F}$ satisfy the Lipschitz condition in the sense that, there exits a constant $M>0$ such that, for all $(x_{i},\mu_{i},y_{i},z_{i})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ with $i=1,2$,
    $|\delta(t,x_{1},\mu_{1},y_{1},z_{1})-\delta(t,x_{2},\mu_{2},y_{2},z_{2})|\leq M(|x_{1}-x_{2}|+\mathcal{W}_{2,\mathbb{R}^{n}}(\mu_{1},\mu_{2})+|y_{1}-y_{2}|+|z_{1}-z_{2}|),$
    for $\delta\in\{\hat{b},\hat{\sigma},\hat{f},\nabla_{x}b,\partial_{\mu}b(\cdot)(x),\nabla_{x}\sigma,\partial_{\mu}\sigma(\cdot)(x),\nabla_{x}f,\partial_{\mu}f(\cdot)(x)\}$ and $x\in\mathbb{R}^{n}$.
Note that Assumption 2 strengthens Assumption 1 by requiring the differentiability of the coefficients and imposing growth conditions on their derivatives, while it is weaker than Assumption 1 by relaxing the growth condition on the running cost function $f$ and omitting the requirement for a uniform positive constant $c_{0}$ of the cost function $c$. Assumption 3 introduces the monotonicity to the coefficients and guarantees the existence of a Lipschitz continuous minimizer for the Hamiltonian. The following flowchart illustrates how the above three assumptions are applied throughout the paper.
Assumption 1Assumption 2Assumption 3Theorem 3.9Corollary 3.10Theorem 4.4Theorem 5.3Theorem 5.5Existence of optimal controlExistence of FBSDEUniqueness of FBSDEStability of FBSDE
@@ 3.1 Relaxed control formulation
To confirm the existence of an optimal pair of regular and singular controls to the problem (3), we first reformulate the problem in a relaxed control framework such that the compactification method can be utilized. To this end, let us introduce the following basic spaces:
  • The space $D^{n}[0,T]$ is endowed with the Skorokhod metric $d_{D}$ and the corresponding Borel $\sigma$-algebra denoted by $\mathcal{F}^{D}$. Let $\mathcal{F}_{t}^{D}$ be the Borel $\sigma$-algebra up to time $t$ (c.f. Haussmann [20]) for $t\in[0,T]$.
  • The space $\mathcal{Q}$ of relaxed controls is defined as the set of measures $q$ in $[0,T]\times U$ with the first marginal equal to the Lebesgue measure and $\int_{[0,T]\times U}|u|^{p}q(\mathrm{d}t,\mathrm{d}u)<\infty$. We endow the space $\mathcal{Q}$ with the $2$-Wasserstein metric on $\mathcal{P}_{2}([0,T]\times U)$ given by $d_{\mathcal{Q}}(q^{1},q^{2})=\mathcal{W}_{2,[0,T]\times U}\left(\frac{q^{1}}{T},\frac{q^{2}}{T}\right)$, where the metric on $[0,T]\times U$ is given by $((t_{1},u_{1}),(t_{2},u_{2}))\mapsto|t_{2}-t_{1}|+|u_{2}-u_{1}|$. Note that each $q\in\mathcal{Q}$ can be identified with a measurable function $[0,T]\in t\mapsto q_{t}\in\mathcal{P}_{2}(U)$, defined uniquely up to $\mathrm{a.s.}$ by $q(\mathrm{d}t,\mathrm{d}u)=q_{t}(\mathrm{d}u)\mathrm{d}t$. In the sequel, we will always refer to the measurable mapping $q=(q_{t})_{t\in[0,T]}$ to a relaxed control in $\mathcal{Q}$. Let $\mathcal{F}^{\mathcal{Q}}$ be the Borel $\sigma$-algebra of $\mathcal{Q}$ and $\mathcal{F}_{t}^{\mathcal{Q}}$ be the $\sigma$-algebra generated by the maps $q\mapsto q([0,s]\times V)$ with $s\in[0,t]$ and Borel measurable $V\subset U$.
  • The space $A^{k}[0,T]$ is a subset of $\mathcal{V}\subset\mathrm{BV}^{k}[0,T]$, where the set $\mathcal{V}\subset\mathrm{BV}^{k}[0,T]$ is the collection of mappings $a:[0,T]\to\mathbb{R}^{k}$ such that $a$ is càdlàg and $a(0)=0$. As in [20], there is a natural one-to-one correspondence between $\mathcal{V}$ and $(\mathcal{M}[0,T])^{k}$, that is, for $a=(a^{1},\ldots,a^{k})\in\mathcal{V}$ and $i=1,\ldots,k$, define $\nu_{i}((s,t])=a^{i}(t)-a^{i}(s)$ for $0\leq s\leq t\leq T$. Then, $\mathcal{V}$ is the dual space of $(C[0,T]^{*})^{k}$, and hence we endow $\mathcal{V}$ with the corresponding weak topology (which is also equivalent to the Fortet-Mourier norm on ($\mathcal{M}[0,T])^{k}$, c.f. [5]). Thus, $A^{k}[0,T]$ is a closed subset of $(\mathcal{M}[0,T])^{k}$ and can also be viewed as $(\mathcal{M}^{+}[0,T])^{k}$. Let $\mathcal{F}^{A}$ be the Borel $\sigma$-algebra and $\mathcal{F}^{A}_{t}$ be the Borel $\sigma$-algebra defined up to time $t\in[0,T]$ (cf. [20]).
@@
Equip $\widehat{\Omega}$ and $\Omega$ with the respective product metrics $d_{\widehat{\Omega}}(\widehat{\omega}^{1},\widehat{\omega}^{2})=d_{D}(\psi^{1},\psi^{2})+d_{\mathcal{Q}}(q^{1},q^{2})$ and $d_{\Omega}(\omega^{1},\omega^{2})=d_{\widehat{\Omega}}(\widehat{\omega}^{1},\widehat{\omega}^{2})+\sum_{i=1}^{k}d_{\rm FM}(v_{i}^{1},\nu_{i}^{2})$ for $\omega^{i}=(\widehat{\omega}^{i},\zeta^{i})=(\psi^{i},q^{i},\zeta^{i})\in\Omega$ with $i=1,2$. Here, $\nu^{i}$ is the correspondence of $\zeta^{i}$ in $\mathcal{V}$ and $d_{\rm FM}$ is the Fortet-Mourier metric on $\mathcal{M}[0,T]$. Consequently, both $(\widehat{\Omega},d_{\widehat{\Omega}})$ and $(\Omega,d_{\Omega})$ are Polish spaces.
Set $\mathcal{F}=\mathcal{F}^{D}\otimes\mathcal{F}^{\mathcal{Q}}\otimes\mathcal{F}^{A}$ as the product $\sigma$-algebra and define the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ by $\mathcal{F}_{t}=\mathcal{F}_{t}^{D}\otimes\mathcal{F}_{t}^{\mathcal{Q}}\otimes\mathcal{F}_{t}^{A}$ for $t\in[0,T]$. For any $\omega=(\psi,q,\zeta)\in\Omega$, Let $X_{t}(\omega)=\psi_{t}$, $q_{t}(\omega)=q_{t}$, and $\zeta_{t}(\omega)=\zeta_{t}$ for $t\in[0,T]$ be the coordinate process. Denote by $\mathcal{P}_{2}(\Omega;\widehat{\Omega})$ the set of probability measures $R$ on $(\Omega,\mathcal{F})$ such that $R|_{\widehat{\Omega}}\in\mathcal{P}_{2}(\widehat{\Omega})$. Introduce the corresponding metric on $\mathcal{P}_{2}(\Omega;\widehat{\Omega})$ by, for all $R_{1},R_{2}\in\mathcal{P}_{2}(\widehat{\Omega})$,
@@
where $d_{\rm FM,\Omega}$ is the Fortet-Mourier metric on $\mathcal{P}(\Omega)$. As a result, $R_{n}\to R$ in $\mathcal{P}_{2}(\Omega;\widehat{\Omega})$ if and only if $R_{n}|_{\widehat{\Omega}}\to R|_{\widehat{\Omega}}$ in $\mathcal{P}_{2}(\widehat{\Omega})$ and $R_{n}$ weakly converges to $R$ as $n\to\infty$.
Next, we give the definition of relaxed controls. In particular, one important task is to guarantee that the dynamic state-control-law constraints are fulfilled when applying the compactification arguments. For this purpose, we propose to embed the constraints into the definition of admissible control rules in the following sense. Recall that $\nu\in\mathcal{P}_{p}(\mathbb{R}^{n})$ is the law of initial state in SDE (1).
In the sequel, we will always consider the weak formulation, $\mathrm{i.e.}$, we will take $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ introduced in last section to be the canonical space.
@@
An admissible control rule is a probability measure $R\in\mathcal{P}_{2}(\Omega;\widehat{\Omega})$ on $(\Omega,\mathcal{F})$ such that
@@
    $M_{t}^{R}g$ $=g(X_{t})-\int_{0}^{t}\int_{U}\mathbb{G}g(s,X_{s},R^{X}(s),u)q_{s}(\mathrm{d}u)\mathrm{d}s-\int_{0}^{t}\nabla_{x}g(X_{s-})G(s)\mathrm{d}\zeta_{s}$
@@
    with $R^{X}(s)=R\circ X_{s}^{-1}$, $\Delta X_{s}=X_{s}-X_{s-}$ and
    $\mathbb{G}g(t,x,\mu,u)=b(t,x,\mu,u)\nabla_{x}g(x)+\frac{1}{2}\mathrm{tr}\left(\sigma\sigma^{\top}(t,x,\mu,u)\nabla_{xx}^{2}g(x)\right).$
@@
    for $i\in I$, $m\times R(A_{i}^{R})=T$ with the admissible set $A_{i}^{R}$ defined by
    $A_{i}^{R}=\left\{(t,\omega)\in[0,T]\times\Omega;~{}\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\geq 0\right\},$
    where $m\times R$ denotes the product measure of Lebesugue measure $m$ and $R$.
@@ 3.2 Existence of optimal relaxed controls
For $\nu\in\mathcal{P}_{p}(\mathbb{R}^{n})$, let $\mathcal{R}_{\rm ad}(\nu)$ be the set of admissible rules $R$ in Definition 3.1 in the previous subsection. For $R\in\mathcal{R}_{\rm ad}(\nu)$, we have the following martingagle measure characterization:
@@
We have $R\in\mathcal{R}_{\rm ad}(\nu)$ if and only if there is a filtered probability space $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{F}^{\prime}=(\mathcal{F}^{\prime}_{t})_{t\in[0,T]},P^{\prime})$ supporting an $n$-dimensional $\mathbb{F}^{\prime}$-adapted process $X=(X_{t})_{t\in[0,T]}$, $r$-dimensional orthogonal $\mathbb{F}^{\prime}$-martinagle measures $N=(N^{1},\ldots,N^{r})$ on $U\times[0,T]$, each with $\mathcal{F}^{\prime}_{t}$-progressively measurable intensity $q_{t}(\mathrm{d}u)\mathrm{d}t$ and a $k$-dimensional $\mathbb{F}^{\prime}$-progressively measurable nondecreasing process $\zeta=(\zeta_{t})_{t\in[0,T]}$ such that $R=P^{\prime}\circ(X,q,\zeta)^{-1}$, and the following conditions hold:
@@
    the state dynamics obeys that, $P^{\prime}$-a.s.,
@@
    the constraints hold that, for $m\times P^{\prime}$-$\mathrm{a.s.}$, $\int_{U}\phi^{i}(t,X_{t},\mathcal{L}^{P^{\prime}}(X_{t}),u)q_{t}(du)\geq 0$ for $i\in I$.
Moreover, there exists a constant $C>0$ depending on $M,\nu,p,T$ such that $\mathbb{E}^{P^{\prime}}[\sup_{t\in[0,T]}|X_{t}|^{p}]\leq C\left\{1+\mathbb{E}\left[|\xi|^{p}\right]\right\}$ with $M$ stated in Assumption 1.
@@
The first assertion follows from El Karoui [16], and the second assertion can be easily proved by basic moment estimations under Assumption 1. ∎
@@
Then, it follows from Assumption 1 that $v^{*}:=\inf_{R\in\mathcal{R}_{\rm ad}(\nu)}\Gamma(R)\in\mathbb{R}$. Consequently, the optimal control rule set is given by
$\mathcal{R}_{\rm opt}(\nu)=\inf\{R\in\mathcal{R}_{\rm ad}(\nu);~{}\Gamma(R)=v^{*}\}.$ (6)
@@
For any strict admissible control $(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]$, if let $X^{\alpha,\zeta}$ be the controlled state process obeyed by (1) and $R$ be the joint law of $(X^{\alpha,\zeta}_{\cdot},\delta_{\alpha_{\cdot}}(\mathrm{d}u)\mathrm{d}t,\zeta_{\cdot})$, then $R$ can be easily verified to be an admissible control rule and $J(\alpha,\zeta)=\Gamma(R)$ with $J(\alpha,\zeta)$ defined in (3). This yields that
$v^{*}\leq\inf_{(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]}J(\alpha,\zeta).$
To arrive at the reverse inequality in Remark 3.3, for any $(t,x,\nu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$, we define a subset of $\mathbb{R}^{n}\times\mathbb{R}^{n\times n}\times\mathbb{R}\times\mathbb{R}^{m}$ as follows:
@@
:=$\qquad:=\left\{(b(t,x,\nu,u),\sigma\sigma^{\top}(t,x,\nu,u),z,w);~{}z\geq f(t,x,\nu,u),w_{i}\leq\phi^{i}(t,x,\nu,u),~{}i\in I,u\in U\right\}.$
@@
Then, we have
@@
For any $\omega\in\Omega$, define the set $K(t,\omega):=K(t,X_{t}(\omega),\mathcal{L}(X_{t}^{\alpha,\zeta}))$ with $(t,\omega)\in[0,T]\times\Omega$. Since $K(t,\omega)$ is convex and closed, we have

### v2

@@ 2.2 Standing assumptions
    Define $F(t,x,\mu,u,y,z):=\nabla_{x}H(t,x,\mu,u,y,z)+\int_{\mathbb{R}^{n}}\partial_{\mu}H(t,x^{\prime},\mu,u,y,z)(x)\mu(\mathrm{d}x^{\prime})$,
    $\hat{F}(t,x,\mu,y,z):=F(t,x,\mu,\hat{u}(t,x,\mu,y,z),y,z)$ and $\Phi(t,x,\mu,y,z):=(-\hat{F},\hat{b},\hat{\sigma})(t,x,\mu,y,z)$. The following monotone condition holds that
@@
    where $\beta>\frac{M}{2}$ with $M$ being given in (A3). The inner product is read as
@@
    $\hat{b}$, $\hat{\sigma}$, $\hat{F}$ are Lipschitz continuous, i.e., there is $M>0$ such that, for $(x_{i},\mu_{i},y_{i},z_{i})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ with $i=1,2$, $\delta\in\{\hat{b},\hat{\sigma},\hat{F}\}$ and $x\in\mathbb{R}^{n}$,
    $|\delta(t,x_{1},\mu_{1},y_{1},z_{1})-\delta(t,x_{2},\mu_{2},y_{2},z_{2})|$
Assumption 2 strengthens Assumption 1 by requiring the differentiability of the coefficients and imposing growth conditions on their derivatives, while it is weaker than Assumption 1 by relaxing the growth condition on the running cost function $f$ and omitting the requirement for a uniform positive constant $c_{0}$ of the cost function $c$. Assumption 3 introduces the monotonicity to the coefficients and guarantees the existence of a Lipschitz continuous minimizer for the Hamiltonian.
@@ 3.1 Relaxed control formulation
To confirm the existence of an optimal pair of regular and singular controls to problem (3), we first reformulate the problem in a relaxed control framework such that the compactification method can be utilized. To this end, let us introduce the following basic spaces:
  • The space $D^{n}[0,T]$ is endowed with the Skorokhod metric $d_{D}$ and the corresponding Borel $\sigma$-algebra denoted by $\mathcal{F}^{D}$. Let $\mathcal{F}_{t}^{D}$ be the Borel $\sigma$-algebra up to time $t$ (c.f. [20]) for $t\in[0,T]$.
  • The space $\mathcal{Q}$ of relaxed controls is defined as the set of measures $q$ in $0,T]\times U$ with the first marginal equal to the Lebesgue measure and $\int_{[0,T]\times U}|u|^{p}q(\mathrm{d}t,\mathrm{d}u)<\infty$. We endow the space $\mathcal{Q}$ with the $2$-Wasserstein metric on $\mathcal{P}_{2}([0,T]\times U)$ given by $d_{\mathcal{Q}}(q^{1},q^{2})=\mathcal{W}_{2,[0,T]\times U}\left(\frac{q^{1}}{T},\frac{q^{2}}{T}\right)$, where the metric on $0,T]\times U$ is given by $((t_{1},u_{1}),(t_{2},u_{2}))\mapsto|t_{2}-t_{1}|+|u_{2}-u_{1}|$. Note that each $q\in\mathcal{Q}$ can be identified with a measurable function $0,T]\in t\mapsto q_{t}\in\mathcal{P}_{2}(U)$, defined uniquely up to $\mathrm{a.s.}$ by $q(\mathrm{d}t,\mathrm{d}u)=q_{t}(\mathrm{d}u)\mathrm{d}t$. In the sequel, we will always refer to the measurable mapping $q=(q_{t})_{t\in[0,T]}$ to a relaxed control in $\mathcal{Q}$. Let $\mathcal{F}^{\mathcal{Q}}$ be the Borel $\sigma$-algebra of $\mathcal{Q}$ and $\mathcal{F}_{t}^{\mathcal{Q}}$ be the $\sigma$-algebra generated by the maps $q\mapsto q([0,s]\times V)$ with $s\in[0,t]$ and Borel measurable $V\subset U$.
  • The space $A^{k}[0,T]$ is a subset of $\mathcal{V}\subset\mathrm{BV}^{k}[0,T]$, where the set $\mathcal{V}\subset\mathrm{BV}^{k}[0,T]$ is the collection of mappings $a:[0,T]\to\mathbb{R}^{k}$ such that $a$ is càdlàg and $a(0)=0$. As in [20], there is a natural one-to-one correspondence between $\mathcal{V}$ and $(\mathcal{M}[0,T])^{k}$, that is, for $a=(a^{1},\ldots,a^{k})\in\mathcal{V}$ and $i=1,\ldots,k$, define $\nu_{i}((s,t])=a^{i}(t)-a^{i}(s)$ for $0\leq s\leq t\leq T$. Then, $\mathcal{V}$ is the dual space of $(C[0,T])^{k}$, and hence we endow $\mathcal{V}$ with the corresponding weak topology (which is also equivalent to the Fortet-Mourier norm on ($\mathcal{M}[0,T])^{k}$, c.f. [5]). Thus, $A^{k}[0,T]$ is a closed subset of $(\mathcal{M}[0,T])^{k}$ and can also be viewed as $(\mathcal{M}^{+}[0,T])^{k}$. Let $\mathcal{F}^{A}$ be the Borel $\sigma$-algebra and $\mathcal{F}^{A}_{t}$ be the Borel $\sigma$-algebra defined up to time $t\in[0,T]$ (cf. [20]).
@@
[…]
@@
[…]
Next, we give the definition of relaxed controls. In particular, one important task is to guarantee that the dynamic state-control-law constraints are fulfilled when applying the compactification argument. For this purpose, we propose to embed the constraints into the definition of admissible control rules in the following sense. Recall that $\nu\in\mathcal{P}_{p}(\mathbb{R}^{n})$ is the law of initial state in SDE (1).
In the sequel, we will always consider the weak formulation, $\mathrm{i.e.}$, we will take $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ introduced in this subsection to be the canonical space.
@@
[…]
@@
    $M_{t}^{R}g$ $:=g(X_{t})-\int_{0}^{t}\int_{U}\mathbb{G}g(s,X_{s},R^{X}(s),u)q_{s}(\mathrm{d}u)\mathrm{d}s-\int_{0}^{t}\nabla_{x}g(X_{s-})G(s)\mathrm{d}\zeta_{s}$
@@
    with $R^{X}(s):=R\circ X_{s}^{-1}$, $\Delta X_{s}:=X_{s}-X_{s-}$ and
    $\mathbb{G}g(t,x,\mu,u):=b(t,x,\mu,u)\nabla_{x}g(x)+\frac{1}{2}\mathrm{tr}\left(\sigma\sigma^{\top}(t,x,\mu,u)\nabla_{xx}^{2}g(x)\right).$
@@
    for $i\in I$, $m\otimes R(A_{i}^{R})=T$ with the admissible set $A_{i}^{R}$ defined by
    $A_{i}^{R}=\left\{(t,\omega)\in[0,T]\times\Omega;~\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\geq 0\right\},$ (5)
    where $m\otimes R$ denotes the product measure of Lebesugue measure $m$ and $R$.
@@ 3.2 Existence of optimal relaxed controls
For $\nu\in\mathcal{P}_{p}(\mathbb{R}^{n})$, let $\mathcal{R}_{\rm ad}(\nu)$ be the set of admissible control rules $R$ in Definition 3.1. For $R\in\mathcal{R}_{\rm ad}(\nu)$, we have the following martingagle measure characterization:
@@
$R\in\mathcal{R}_{\rm ad}(\nu)$ iff there is a filtered probability space $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{F}^{\prime}=(\mathcal{F}^{\prime}_{t})_{t\in[0,T]},P^{\prime})$ supporting an $n$-dimensional $\mathbb{F}^{\prime}$-adapted process $X=(X_{t})_{t\in[0,T]}$, $r$-dimensional orthogonal $\mathbb{F}^{\prime}$-martingale measures $N=(N^{1},\ldots,N^{r})$ on $U\times[0,T]$, each with $\mathcal{F}^{\prime}_{t}$-progressively measurable intensity $q_{t}(\mathrm{d}u)\mathrm{d}t$ and a $k$-dimensional $\mathbb{F}^{\prime}$-progressively measurable nondecreasing process $\zeta=(\zeta_{t})_{t\in[0,T]}$ such that $R=P^{\prime}\circ(X,q,\zeta)^{-1}$, and the following conditions hold:
@@
    the state dynamics obeys that, $P^{\prime}$-a.s.
@@
    the constraints hold that, for $m\otimes P^{\prime}$-$\mathrm{a.s.}$, $\int_{U}\phi^{i}(t,X_{t},\mathcal{L}^{P^{\prime}}(X_{t}),u)q_{t}(du)\geq 0$ for $i\in I$.
Moreover, there exists $C>0$ depending on $M,\nu,p,T$ such that $\mathbb{E}^{P^{\prime}}[\sup_{t\in[0,T]}|X_{t}|^{p}]\leq C\left\{1+\mathbb{E}\left[|\xi|^{p}\right]\right\}$ with $M$ stated in Assumption 1.
@@
The first assertion follows from [16], and the second assertion can be easily proved by basic moment estimations under Assumption 1. ∎
@@
By using Assumption 1, $v^{*}:=\inf_{R\in\mathcal{R}_{\rm ad}(\nu)}\Gamma(R)\in\mathbb{R}$. Then, the optimal control rule set is given by
$\mathcal{R}_{\rm opt}(\nu):=\inf\{R\in\mathcal{R}_{\rm ad}(\nu);~\Gamma(R)=v^{*}\}.$ (7)
@@
For any strict admissible control $(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]$, let $X^{\alpha,\zeta}$ be the state process obeyed by (1) and $R$ be the joint law of $(X^{\alpha,\zeta}_{\cdot},\delta_{\alpha_{\cdot}}(\mathrm{d}u)\mathrm{d}t,\zeta_{\cdot})$. Then, $R$ can be easily verified to be an admissible control and $J(\alpha,\zeta)=\Gamma(R)$ with $J(\alpha,\zeta)$ defined in (3). This yields that $v^{*}\leq\inf_{(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]}J(\alpha,\zeta)$.
To arrive at the reverse inequality in Remark 3.3, for $(t,x,\nu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$, define a subset of $\mathbb{R}^{n}\times\mathbb{R}^{n\times n}\times\mathbb{R}\times\mathbb{R}^{m}$ by
@@
$\quad:=\left\{(b(t,x,\nu,u),\sigma\sigma^{\top}(t,x,\nu,u),z,w);~z\geq f(t,x,\nu,u),w_{i}\leq\phi^{i}(t,x,\nu,u),~i\in I,u\in U\right\}.$
@@
Then, we have the following lemma whose proof is provided in Appendix.
@@
We next study the existence of optimal relaxed controls. Consider the McKean-Vlasov SDE (1) and define $Y_{t}:=X_{t}^{\alpha,\zeta}-\int_{0}^{t}G(s)\mathrm{d}\zeta_{s}$ for $t\in[0,T]$. It follows from the dynamics (1) that
Motivated by this expression, we have the next result whose proof is omitted because the necessary part is obvious and the sufficiency can be shown by the classical martingale argument.
Lemma 3.5.
[…]
  • (i)
    $Y$ has continuous trajectories and $R(X_{\cdot}=Y_{\cdot}+\int_{0}^{\cdot}G(s)\mathrm{d}\zeta_{s})=1$; (ii) $R\circ Y_{0}^{-1}=\nu$;
  • (iii)
    $\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}|u|^{p}q_{t}(\mathrm{d}u)\mathrm{d}t+|\zeta_{T}|^{p}\right]<\infty$; (iv) for any $g\in C_{b}^{2}(\mathbb{R}^{n})$, $\tilde{M}_{t}^{R}g$ is a $(R,\mathbb{F})$-martingale, where $\tilde{M}_{t}^{R}g:=g(Y_{t})-\int_{0}^{t}\int_{U}\tilde{\mathbb{G}}g(s,X_{s},R^{X}(s),u)(Y_{s})q_{s}(\mathrm{d}u)\mathrm{d}s$ with
    $\tilde{\mathbb{G}}g(t,x,\mu,u)(y):=b(t,x,\mu,u)\nabla_{x}g(y)+\frac{1}{2}\mathrm{tr}(\sigma\sigma^{\top}(t,x,\mu,u)\nabla_{xx}^{2}g(y)).$
  • (v)
    for $i\in I$, we have $m\otimes R(A_{i}^{R})=T$.
Moreover, we have
Lemma 3.6.
The objective functional $R\mapsto\Gamma(R)$ defined by (6) is l.s.c. on $\mathcal{R}_{\rm ad}(\nu)$.